The principal of tala senior highschool claims that the students in his school have above average intellegence. A random sample of 30 students IQ scores have a mean score of 113. The mean population IQ is 100 with a standard deviation of 15. Is there an evidence to support his claim?


1
Expert's answer
2022-05-05T11:05:31-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\le100"

"H_a:\mu>100"

This corresponds to a right-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a right-tailed test is "z_c = 1.6449."

The rejection region for this right-tailed test is "R = \{z: z > 1.6449\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{113-100}{15/\sqrt{30}}\approx4.7469"

Since it is observed that "z = 4.7469 > 1.6449=z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=P(Z>4.7469)= 0.000001," and since "p= 0.000001<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is greater than 100, at the"\alpha = 0.05" significance level.